OpenStudy (anonymous):

Pls help me solve, thanks Find the distance d(P1 and P2) between the points P1 and P2 P1 = (-2, 5) P2 = (1,4)

OpenStudy (anonymous):

\[d=\sqrt{(x2-x1)^2+(y2-y1)^2}\]

OpenStudy (anonymous):

Thanks. Not good at square roots but this will do. Thanks

OpenStudy (anonymous):

is the ans the square root of -5?

OpenStudy (anonymous):

that is 2.24

OpenStudy (anonymous):

I got \[\sqrt{10}\]

OpenStudy (anonymous):

uhmm will try it again but thanks.

OpenStudy (anonymous):

1-(-2) = 3; 3^2 = 9 4-5 = -; -1^2 = 1 sqrt(9+1) = sqrt(10)

OpenStudy (anonymous):

*4-5= -1

OpenStudy (anonymous):

Lets say you use 2 as x2 and 1 as x1...the same for the y's. (You should get the same thing)

OpenStudy (anonymous):

ok.... trying again

OpenStudy (anonymous):

-2-1 = -3; -3^2 = 9 5-4 =1 1^2= 1 sqrt(9+1) = sqrt(10) You should get sqrt(10) either way

OpenStudy (anonymous):

Don't sqaure right away, do subtraction first

OpenStudy (anonymous):

I did this time thanks!!

OpenStudy (anonymous):

got it. :)

OpenStudy (anonymous):

I think you were doing sqrt(1- (-2)^2); resulting in sqrt(1-4) instead of sqrt(-3^2)

OpenStudy (anonymous):

that was what i did. you are right

OpenStudy (anonymous):